# CoDet-M4: Detecting Machine-Generated Code in Multi-Lingual, Multi-Generator and Multi-Domain Settings

### v1

@@ Abstract
Large language models (LLMs) have revolutionized code generation, automating programming with remarkable efficiency. However, these advancements challenge programming skills, ethics, and assessment integrity, making the detection of LLM-generated code essential for maintaining accountability and standards. While, there has been some research on this problem, it generally lacks domain coverage and robustness, and only covers a small number of programming languages. To this end, we propose a framework capable of distinguishing between human- and LLM-written code across multiple programming languages, code generators, and domains. We use a large-scale dataset from renowned platforms and LLM-based code generators, alongside applying rigorous data quality checks, feature engineering, and comparative analysis using evaluation of traditional machine learning models, pre-trained language models (PLMs), and LLMs for code detection. We perform an evaluation on out-of-domain scenarios, such as detecting the authorship and hybrid authorship of generated code and generalizing to unseen models, domains, and programming languages. Moreover, our extensive experiments show that our framework effectively distinguishes human- from LLM-written code and sets a new benchmark for this task.
@@ 1 Introduction
Recent advancements in LLMs have demonstrated their remarkable ability to generate outputs that closely mimic human-written content Jiang et al. (2024). This has spurred exponential growth in research, with publications on LLM-based code generation in leading venues rising from a single article in 2018 to 140 in 2024 Chang et al. (2024). Moreover, these developments promise to accelerate software development, automate routine tasks, and boost productivity.
@@
However, rapid progress in artificial intelligence (AI) generative systems has also introduced significant challenges, particularly in ethical use and accountability Al-kfairy et al. (2024). Machine-generated code can be exploited to create obfuscated scripts, introduce vulnerabilities, and produce deceptive artifacts that are difficult to trace Bukhari (2024). Thus, tracing AI-assisted commits empowers the code reviewers to proactively mitigate these risks. In academia, the reliance on LLMs for assignments undermines educational integrity, with professors unknowingly grading machine-generated submissions Koike et al. (2024); Ma et al. (2023). Alarmingly, more than 60,000 scientific articles in the past year alone have shown evidence of machine-generated content Gray (2024). This underscores the urgent need for accountability among both students and educators to uphold educational integrity.
@@
Accountability is equally critical in hiring and evaluation. Employers must verify that the code submitted truly reflects the candidate’s abilities rather than relying on AI assistance. Without robust detection mechanisms, generative AI could lead to misleading assessments. Moreover, distinguishing LLM-generated code is essential for developing more effective code-based LLMs; since these systems depend on human-written samples for training, accurately detecting AI-generated code helps curate higher-quality training datasets.
@@
For this purpose, prior studies have proposed frameworks for detecting machine-generated code, including contrastive learning with a UniXcoder-based semantic encoder Xu et al. (2024a) and ML models analyzing Claude 3-generated code in the CodeSearchNet dataset Rahman et al. (2024). However, these approaches often overlook the growing prevalence of open-source and locally deployable LLM-based code assistants.
@@
Despite previous efforts, this evolving state of LLMs de facto highlights the urgent need for reliable curated large-scale high-quality code data with various programming languages and methods to distinguish between human-written and LLM-generated code to mitigate risks and maintain accountability in software development and academic integrity Chang et al. (2024). With this in mind, addressing the aforementioned limitations is crucial, as LLMs are designed to generate code across various languages and domains [^1] Domain in our study corresponds to combination of both data source (GitHub, LeetCode, Codeforces) and its form: functions, classes, arbitrary code snippets..
@@
Here, we aim to close these gaps by constructing first-of-its-kind a large-scale, multi-lingual, multi-domain, and multi-generator dataset comprising $\approx 500K$ samples of human- and LLM-written code. The dataset spans class-level and function-level code, as well as competitive and industrial programming contexts. We further propose models for detecting LLM-generated code and authorship and hybrid authorship identification, and we evaluate their performance in extreme out-of-domain (OOD) settings, such as detecting unseen models, unseen domains, and unseen programming languages.
We aim to answer the following research questions (RQs):
@@
  • (RQ3) Can detection models generalize robustly across different code-generating models, domains, programming languages and hybrid-authorship scenarios?
Our study makes the following contributions:
  • We introduce a novel corpus and benchmark designed for studying machine-generated and human-written code. Spanning a wide array of models, domains, and programming languages, providing a diverse and large-scale foundational resource for research in this area.
  • We repurpose existing code-related models, fine-tuning them to identify machine-, human- and hybrid-written code. Our analysis identifies the strengths and limitations of these models and offers valuable insight into their performance and applicability.
  • We analyze the performance of these models from multiple perspectives: (i) authorship and hybrid authorship identification, examining how well detectors differentiate the author of the code; (ii) unseen models, evaluating their robustness against previously unencountered code generators; (iii) unseen domains, assessing their adaptability to new programming contexts; (iv) unseen languages, testing their generalization to programming languages not included in the training data; and (v) mixed authorship scenario: when LLMs not just generate the code, but rather complement or rewrite human-written code. These experiments yield key insights to guide and inform future research in this area.
  • We make openly available our data and code [^2] https://huggingface.co/datasets/DaniilOr/CoDET-M4, and we are committed to continuously updating our repository with additional generators, domains, and languages in the future.
@@ 2 Related Work
In this section, we discuss resources and frameworks for detecting machine-generated code and consider machine-generated text frameworks, which provide relevant insights and inspiration for our work.
@@
Resources: Significant advancements have been made in developing benchmarks to evaluate LLMs in code generation, covering various domains. Chen et al. (2021) introduced 164 Python problems with function signatures and unit tests, extended by Liu et al. (2024) with 80x more test cases, while Muennighoff et al. (2023) added tasks such as code synthesis in six languages. Austin et al. (2021a) provided 974 entry-level Python problems and Yin et al. (2018) curated 597K Python code generation samples. Yu et al. (2018) evaluated text-to-SQL queries across 138 domains, Iyer et al. (2018) tested near zero-shot Java class generation, and Wang et al. (2022b) focused on execution-based Python code generation. The generation of pragmatic codes was assessed by Yu et al. (2024), robustness was evaluated by Wang et al. (2022a), while McLean Babe et al. (2023) examined student-authored prompts. Finally, Zhuo et al. (2024), Du et al. (2024), and Zhang et al. (2024) targeted cross-domain, class-level, and multi-lingual tasks, respectively. Athiwaratkun et al. (2022) adapted Austin et al. (2021a) for multiple languages and Zheng et al. (2023) extended for multilingual tasks, evaluating code in C++, Java, JavaScript and Go. Cassano et al. (2022) benchmarked code generation in 18 languages and Khan et al. (2023) provided 25M multilingual examples for multitask evaluations. Austin et al. (2021a) synthesized code from complex descriptions, while Gu et al. (2024) evaluated reasoning and execution capabilities using 800 Python functions. Hendrycks et al. (2021) included 10K Python problems at varying difficulty levels, Li et al. (2022) offered competitive problems with test cases from platforms such as CodeForces, and Jain et al. (2024) evaluated code generation, repair, and execution across 713 coding problems. Chandel et al. (2022) evaluated pedagogical data science notebooks, Lai et al. (2023) introduced 1K science questions covering Python libraries, and Huang et al. (2022) focused on execution-based evaluation using 534 Jupyter Notebook problems.
@@
Machine-generated code detection: Nguyen et al. (2024) proposed a binary classifier to detect ChatGPT-generated code in Python and Java, using the CodeSearchNet dataset Husain et al. (2019). Xu et al. (2024b) demonstrated contrastive learning on a 500K sample parallel corpus to improve detection, and Idialu et al. (2024) employed stylometric features to identify GPT-4-generated code at the class level. However, these studies are limited to function-level or Python-based detection, underscoring the need for broader datasets and methods for diverse languages and domains.
@@
Machine-generated text detection: Wang et al. (2024b) and Guo et al. (2023) created large-scale datasets to improve detection in domains and languages. Statistical methods such as perplexity analysis were introduced by Gehrmann et al. (2019), while Verma et al. (2024) explored text statistics for effective detection. Mitchell et al. (2023) and Bao et al. (2024) showcased tools such as GPTZero and Fast-DetectGPT to distinguish human- and machine-written text, but Pan et al. (2024) revealed limitations in detecting LLM-generated code, emphasizing the need for better solutions.
@@ 3.1 Data Collection
Our work focuses on the most wide-spread programming languages [^3] Python, Java and C++ together account for 1/3 of all pushes, and PRs on github.. We combined data from multiple sources to build our dataset. As a foundation, we used the dataset by Pan et al. (2024), which primarily includes Python code from LeetCode444www.leetcode.com, GeeksForGeeks555www.geeksforgeeks.org, and W3Resource666www.w3resource.com, comprising 5,069 problems with 13 prompts for code generation (dataset under CC BY 4.0 License). Additionally, we collected 2,800 human-written solutions in C++ and Java from LeetCode, we further refer to it as a LeetCode data, focusing on class-level human- and machine-generated code examples. We also retrieved human-written solutions from a publicly available Kaggle dataset777www.kaggle.com, containing 2,523 CodeForces problems with solutions in Python, C++ and Java. Filtering for solutions that passed all CodeForces test cases, this dataset resulted in 103,792 codes: 41,402 in C++, 32,145 in Python and 30,245 in Java.
@@
Figure 1: AST depth comparison.
@@ 3.2 Code Generation
For code generation, we use open-source state-of-the-art models that are lightweight enough (7-8B parameters) to be run locally, aligning with our focus on easily deployable systems. In addition, we included GPT-4o, one of the most accurate and widely used proprietary LLMs, to benchmark against open-source alternatives. To select the most promising open-source models, we refer to the BigCode models leaderboard888www.huggingface.co, which leads us to choose the following: CodeLlama (7B) Rozière et al. (2024), Llama3.1 (8B) Grattafiori et al. (2024), CodeQwen 1.5 (7B) Bai et al. (2023), and Nxcode-orpo (7B), a version of CodeQwen fine-tuned using monolithic preference optimization without reference models Hong et al. (2024).
The generation process employs domain-specific prompts, as shown in the Appendix F. All models were served using vLLM999www.github.com to simulate real-world inference scenarios. To introduce variability in the generated outputs, we used random temperature values ranging from 0.4 to 1.
@@ 3.3 Quality Assurance
Table 2: Evaluation results for different models. The best results are shown in bold.
@@ 3.4 Resulting Dataset
4 Experiments & Results
In this section, we detail our experiments aimed at developing models to detect LLM-generated code. We evaluate these models under extreme conditions, including unseen models, unseen languages, and code from the unseen domains (more precisely: unseen code sources and unseen code structures). Our experimental setup and evaluation measures can be found in Appendix C.
@@ 4 Experiments & Results
Table 3: Evaluation results for models across the three programming languages.
4.1 LLM-generated Code Detection
Regarding RQ1, Table 2 shows that the models can almost perfectly identify LLM-generated code. Even simpler models such as SVM and CatBoost perform considerably better. In Appendix H, we explore what enables these simple models to identify LLM-generated code. Moreover, we also analyze model performance for each programming language, data source, and generator.
Table 4 indicates that the performance of the model varies between different data sources. All models except UniXcoder perform worst on LeetCode data, which could be attributed to mixing LeetCode with other platforms in this set, leading to slight differences in question types. Moreover, confusion matrices for the best model across languages and sources are available in Figure 12, 13 (Appendix I), respectively.
@@ 4.1 LLM-generated Code Detection
Table 4: Evaluation results across the sources.
4.2 Can LLMs Detect Machine-Generated Code and Authorship?
Overall, GPT-4o proved to be ineffective at identifying generated code, even with handcrafted instructions and few-shot samples, leading us to exclude it from further experiments. Handcrafted prompts are available in the Appendix D.
4.3 Authorship Identification
To validate RQ2, which aims to identify the specific model responsible for generating a given piece of code, we conduct experiments using the same experimental setup as described in Appendix C. However, we modified the classification objective: instead of performing binary classification (human-written vs. LLM-generated code), the models are tasked with a multi-class classification problem. This setup involved six distinct classes, representing five different LLMs and human authors.

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized code generation, automating programming with remarkable efficiency. However, this has had important consequences for programming skills, ethics, and assessment integrity, thus making the detection of LLM-generated code essential for maintaining accountability and standards. While, there has been some previous research on this problem, it generally lacks domain coverage and robustness, and only covers a small number of programming languages. Here, we aim to bridge this gap. In particular, we propose a framework capable of distinguishing between human-written and LLM-generated program code across multiple programming languages, code generators, and domains. We use a large-scale dataset from renowned platforms and LLM-based code generators, alongside applying rigorous data quality checks, feature engineering, and comparative analysis of traditional machine learning models, pre-trained language models (PLMs), and LLMs for code detection. We perform an evaluation on out-of-domain scenarios, such as detecting authorship and hybrid authorship of generated code and generalizing to unseen models, domains, and programming languages. Our extensive experiments show that our framework effectively distinguishes human-written from LLM-generated program code, setting a new benchmark for the task.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have demonstrated their remarkable ability to generate outputs that closely emulate human-written content Jiang et al. (2024). This has spurred exponential growth in research, with publications on LLM-based code generation in leading venues rising from a single article in 2018 to 140 in 2024 Chang et al. (2024). Moreover, these developments promise to accelerate software development, automate routine tasks, and boost productivity.
@@
At the same time, the rapid progress in artificial intelligence (AI) generative systems has posed major concerns, particularly related to accountability and ethical use of this technology Al-kfairy et al. (2024). Machine-generated code can be exploited to create obfuscated scripts, introduce vulnerabilities, and produce deceptive artifacts that are difficult to trace Bukhari (2024). Thus, it is important to develop tools that can detect machine-generated content, e.g., tracing AI-assisted commits could empower code reviewers to proactively mitigate risks. In academia, the use of LLMs for completing written assignments undermines educational integrity, with professors unknowingly grading machine-generated submissions Koike et al. (2024); Ma et al. (2023). Alarmingly, more than 60,000 scientific articles in the past year alone have shown evidence of machine-generated content Gray (2024).
@@
Accountability is equally critical for talent hiring and candidate evaluation. Employers must verify that the code submitted by a candidate truly reflects their actual abilities. Without robust detection mechanisms, generative AI could lead to misleading assessments. Moreover, detecting the use of LLM-generated code is essential for developing more effective code-based LLMs; since these systems depend on human-written samples for training, accurately detecting AI-generated code helps curate higher-quality training datasets.
@@
Previous work has proposed frameworks for detecting machine-generated code, including contrastive learning with a UniXcoder-based semantic encoder Xu et al. (2025) and machine learning (ML) models that analyzed Claude 3-generated code in the CodeSearchNet dataset Rahman et al. (2024). That work has focused on a single API-based code copilot, but today we face a growing prevalence of a variety of open-weights LLMs as well as locally deployable LLM-based code assistants.
@@
Despite previous efforts, this evolving state of LLMs de facto highlights the urgent need for reliable curated large-scale high-quality code data with various programming languages and methods to distinguish between human-written and LLM-generated code to mitigate risks and maintain accountability in software development and academic integrity Chang et al. (2024). With this in mind, addressing the aforementioned limitations is crucial, as LLMs are designed to generate code across various languages and domains [^1] Domain in our study corresponds to a combination of data source (GitHub, LeetCode, Codeforces) and its form: functions, classes, and arbitrary code snippets.
@@
Here, we aim to bridge these gaps by constructing a first-of-its-kind large-scale, multi-lingual, multi-domain, and multi-generator dataset comprising $\approx 500K$ samples of human- and LLM-written code. The dataset spans class-level and function-level code, as well as competitive and industrial programming contexts. We further propose models for detecting LLM-generated code, and we evaluate their performance in extreme out-of-domain (OOD) settings: detecting unseen models, unseen domains, and unseen programming languages. We aim to answer the following research questions (RQs):
@@
  • (RQ3) Can detection models generalize robustly across different code-generating models, domains, programming languages, and hybrid-authorship scenarios?
We make the following contributions:
  • We introduce a novel corpus and benchmark designed for studying machine-generated and human-written code. Spanning a wide array of models, domains, and programming languages, providing a diverse and large-scale foundational resource for further research.
  • We repurpose existing code-related models, fine-tuning them to identify machine-, human- and hybrid-written code.
  • We analyze the performance of these models from multiple perspectives: (i) authorship and hybrid authorship identification, (ii) unseen code generators, (iii) unseen domains, (iv) unseen programming languages, and (iv) mixed authorship scenario: when LLMs not just generate the code, but rather complement or rewrite human-written code.
  • We release our data and code [^2] https://huggingface.co/datasets/DaniilOr/oDET-M4, and we are committed to continuously updating our repository with additional generators, domains, and languages in the future.
@@ 2 Related Work
In this section, we discuss related work, focusing on resources and models for detecting machine-generated code.
@@
Resources: Major advancements have been made in developing benchmarks to evaluate LLMs for code generation, covering various domains. Chen et al. (2021) introduced 164 Python problems with function signatures and unit tests, extended by Liu et al. (2023) with 80x more test cases, while Muennighoff et al. (2024) added tasks such as code synthesis in six languages. Austin et al. (2021) provided 974 entry-level Python problems and Yin et al. (2018) curated 597K Python code generation samples. Yu et al. (2018) evaluated text-to-SQL queries across 138 domains, Iyer et al. (2018) tested near zero-shot Java class generation, and Wang et al. (2023c) focused on execution-based Python code generation. The generation of pragmatic codes was assessed by Yu et al. (2024), robustness was evaluated by Wang et al. (2023a), while Babe et al. (2024) examined student-authored prompts. Finally, Zhuo et al. (2025), Du et al. (2024), and Zhang et al. (2024) targeted cross-domain, class-level, and multi-lingual tasks, respectively. Athiwaratkun et al. (2023) adapted Austin et al. (2021) for multiple languages and Zheng et al. (2023) extended for multilingual tasks, evaluating code in C++, Java, JavaScript and Go. Cassano et al. (2023) benchmarked code generation in 18 languages and Khan et al. (2023) provided 25M multilingual examples for multitask evaluations. Austin et al. (2021) synthesized code from complex descriptions, while Gu et al. (2024) evaluated reasoning and execution capabilities using 800 Python functions.
Hendrycks et al. (2021) included 10K Python problems at varying difficulty levels, Li et al. (2022) offered competitive problems with test cases from platforms such as CodeForces, and Jain et al. (2024) evaluated code generation, repair, and execution across 713 coding problems. Chandel et al. (2022) evaluated pedagogical data science notebooks, Lai et al. (2023) introduced 1K science questions covering Python libraries, and Huang et al. (2022) focused on execution-based evaluation using 534 Jupyter Notebook problems.
@@
Machine-generated code detection: Nguyen et al. (2024) proposed a binary classifier to detect ChatGPT-generated code in Python and Java, using the CodeSearchNet dataset Husain et al. (2019). Xu et al. (2025) demonstrated contrastive learning on a 500K sample parallel corpus to improve detection, and Idialu et al. (2024) employed stylometric features to identify GPT-4-generated code at the class level. However, these studies are limited to function-level or Python-based detection, underscoring the need for broader datasets and methods for diverse languages and domains.
@@
Machine-generated text detection: Wang et al. (2024b) and Guo et al. (2023) created large-scale datasets to improve detection across domains, languages, and generators. Abassy et al. (2024) introduced a tool for more fine-grained detection. Statistical methods such as perplexity analysis were introduced by Gehrmann et al. (2019), while Verma et al. (2024) explored text statistics for effective detection. Mitchell et al. (2023) and Bao et al. (2024) showcased tools such as GPTZero and Fast-DetectGPT to distinguish human-written and machine-generated text, but Pan et al. (2024) revealed limitations in the detection of LLM-generated code, emphasizing the need for better solutions.
@@ 3.1 Data Collection
Our work focuses on the most wide-spread programming languages [^3] Python, Java and C++ together account for 1/3 of all pushes, and PRs on github.. We combined data from multiple sources to build our dataset. As a foundation, we used the dataset by Pan et al. (2024), which primarily includes Python code from LeetCode444www.leetcode.com, GeeksForGeeks555www.geeksforgeeks.org, and W3Resource666www.w3resource.com, comprising 5,069 problems with 13 prompts for code generation (dataset under CC BY 4.0 License). Additionally, we collected 2,800 human-written solutions in C++ and Java from LeetCode, we further refer to it as a LeetCode data, focusing on class-level human- and machine-generated code examples. We also retrieved human-written solutions from a publicly available Kaggle dataset777www.kaggle.com, containing 2,523 CodeForces problems with solutions in Python, C++, and Java. Filtering for solutions that passed all CodeForces test cases, this dataset resulted in 103,792 codes: 41,402 in C++, 32,145 in Python and 30,245 in Java.
@@
Figure 1: AST depth comparison between human- and AI-authored codes.
@@ 3.2 Code Generation
For code generation, we use open-source state-of-the-art models that are lightweight enough (7-8B parameters) to be run locally, aligning with our focus on easily deployable systems. In addition, we included GPT-4o, one of the most accurate and widely used proprietary LLMs, to benchmark against open-source alternatives. To select the most promising open-source models, we refer to the BigCode models leaderboard888www.huggingface.co, which leads us to choose the following: CodeLlama (7B) Codellama (2023), Llama3.1 (8B) Llama (2024), CodeQwen 1.5 (7B) Qwen (2023), and Nxcode-orpo (7B), a version of CodeQwen fine-tuned using monolithic preference optimization without reference models Hong et al. (2024).
The generation process employs domain-specific prompts, as shown in the Appendix E. All models were served using vLLM999www.github.com to simulate real-world inference scenarios. To introduce variability in the generated outputs, we used random temperature values ranging from 0.4 to 1.
@@
Table 2: Binary classification results for different models. The best results are shown in bold. P: precision, R: recall, F: F1-score, A: accuracy.
3.3 Quality Assurance
[…]
@@ 3.4 Resulting Dataset
Table 3: Binary classification results for models across the three programming languages.
4 Experiments & Results
In this section, we detail our experiments aimed at developing models to detect LLM-generated code. We evaluate these models under extreme conditions, including unseen models, unseen languages, and code from the unseen domains (more precisely: unseen code sources and unseen code structures).
4.1 Experimental Setup
For the traditional approach, we followed a methodology similar to Idialu et al. (2024), using SVM and the CatBoost gradient booster algorithm Prokhorenkova et al. (2018) to make predictions based on the statistical features of the code. These features included average line length, maximum length of decision operators, function density (number of function definitions per line of code), average function length, whitespace ratio, average variable name length, maintainability index, Abstract Syntax Trees (AST) depth, number of assignment operators, and AST node density for all AST node types. This resulted in over 500 features. Since not all code samples shared the same properties, many features were sparse. To address this, we retained only the features with no more than 20% missing values. Given that the number of features was significantly smaller than the number of samples, we trained the SVM with an RBF kernel using the primal formulation instead of the dual. For the CatBoost model, we trained 2,000 trees, as determined to be optimal based on a grid search optimizing the validation F1-score. Additionally, the learning rate for CatBoost was automatically set to 0.1, which balanced convergence speed and performance.
For DNN-based methods, we tested multiple models that serve as code encoders. CodeBERT, a variant of the BERT model pre-trained on both text and code data Feng et al. (2020). UniXcoder, a model with cross-modal (AST and text) representation of text and code, trained to be used as encoder, decoder, or both Guo et al. (2022). CodeT5, a T5 fine-tuning for multiple code-related tasks such as code completion, text-to-code generation, code retrieval, duplicate detection, etc. Wang et al. (2023b). All of these models were trained in similar settings: for five epochs with initial learning rate of $43e-$ 4, weight decay of $31e-$ 3, batch size of 256, and a linear learning rate scheduler.
@@ 4.1 Experimental Setup
Table 4: Binary classification results across the sources.
Evaluation Measures: To evaluate the performance of the models, we used the Macro F1 score [^10] It balances importance of all classes. (F), precision (P), and recall (R). We also report accuracy (A), since the classes are nearly balanced.
4.2 LLM-generated Code Detection
Regarding RQ1, Table 2 shows that the models can almost perfectly identify the LLM-generated code. Even simpler models such as SVM and CatBoost perform considerably better. In Appendix G, we explore what enables these simple models to identify LLM-generated code. Moreover, we also analyze the performance of the model for each programming language, data source, and generator.
Table 4 indicates that the performance of the model varies between different data sources. All models except UniXcoder perform worst on LeetCode data, which could be attributed to mixing LeetCode with other platforms in this set, leading to slight differences in question types. Moreover, confusion matrices for the best model across languages and sources are available in Figures 12 and 13 (Appendix H), respectively.
@@ 4.2 LLM-generated Code Detection
4.3 Can LLMs Detect Machine-Generated Code and Authorship?
Overall, GPT-4o proved to be ineffective at identifying generated code, even with handcrafted instructions and few-shot samples, leading us to exclude it from further experiments. Handcrafted prompts are available in the Appendix C.
4.4 Authorship Identification
To validate RQ2, which aims to identify the specific model responsible for generating a given piece of code, we conduct experiments using the same experimental setup as described in the § 4.1. However, we modified the classification objective: instead of performing binary classification (human-written vs. LLM-generated code), the models are tasked with a multi-class classification problem. This setup involved six distinct classes, representing five different LLMs and human authors.